Find dy/dx for the equation y= 3x-5/×^2-4x

1
Expert's answer
2022-03-20T06:42:50-0400

"\frac{dy}{dx}=\frac{d}{dx}\left( 3x-\frac{5}{x^2-4x} \right) =3-5\left( -\frac{1}{\left( x^2-4x \right) ^2}\cdot \left( 2x-4 \right) \right) =\\=3+\frac{10\left( x-2 \right)}{\left( x^2-4x \right) ^2}"